# Kibble-Zurek behavior in a topological phase transition with a quadratic band crossing

### v1

@@ VI Summary
In this work, we investigated the KZ behavior in a topological phase transition with a quadratic band crossing. We used dual perspectives: momentum space in clean systems and real space in disordered systems. Both perspectives give consistent scalings, which is distinct from previously studied linear band crossing models. Further, we go beyond quadratic to higher orders. They also follow KZ behavior but acquire different critical exponents.
Our study extends the KZ behavior in topological transitions beyond linear band crossing and confirms the consistency between correlation length from Berry curvature and LCM. For future studies, we get some hint from this work that the LCM correlation length may characterize the Anderson insulator transition if the disorder is strong.
@@ Appendix A Extraction of correlation length from the local Chern marker
Figure 8: (a). The distribution of the LCM at freeze-out time for different quench time. (b). Disorder-averaged autocovariance function calculated from the LCM. The inset shows the same data in a non-rescaled coordinate. Here, the NNN hopping strength is set to $J^{\prime}=1$, and the average is performed over $10$ disorder realizations $\delta V=0.05$ and size $N=50$.
Fig.8(a) shows the real-space distribution of the local Chern marker (LCM) at freeze-out time for different quench time during the quench process. To estimate the size of inhomogeneities in the LCM, we calculated the normalized autocovariance function of the real-space distribution function $\delta c(r)$:
@@
where $\delta c(\mathbf{r})=c(\mathbf{r})-\bar{c}$ is the LCM fluctuation. Similar to previous work, we determine the correlation length in the LCM as the distance at which the disorder-averaged autocovariance function crosses zero, $\bar{R}(\xi_{r})=0$. The autocovariance function of the LCM at freeze-out time for different quench time, is shown in Fig.8(b), plotted as a function of $r/\tau^{1/4}$. Since the correlation length at the freeze-out time satisfies $\xi_{r}\sim\tau^{1/4}$, The autocovariance functions of the LCM in Fig.8(b) overlap near $\bar{R}(\xi_{r})=0$.
@@ Appendix B Alternative characterization of the freeze-out time in momentum space
Figure 9: (a). The dynamics of correlation length in momentum space in a rescaled coordinate. Curves collapse near the transition point. The freeze-out time $t_{\mathrm{f}}$ can be determined by the first peak after the transition. The inset shows the same data in a non-rescaled coordinate. (b). Freeze-out time $t_{\mathrm{f}}$ has a power-law scaling with the quench time $t_{\mathrm{f}}\sim\tau^{1/2}$. (c). Correlation length at freeze-out time $t_{\mathrm{f}}$ shows a scaling behavior $\xi(t_{\mathrm{f}})\sim\tau^{1/4}$.
In the main text, we use the nonequilibrium Hall response to determine the freeze-out time during the quench in clean systems. The freeze-out time by definition characterizes the time when the system’s relaxation rate matches the cooling rate. That is to say, it could be seen as a time delay when the system realizes the transition has occurred. So, the time delay of the trend of divergence of the correlation length can also be utilized to identify the freeze-out time $t_{\mathrm{f}}$. This method can be regarded as a supplement to the Hall response criterion and can be generalized to real space for disordered systems directly, while Hall response is not easy to find analog in real space.

### v2

@@ VI Summary
In this work, we investigated the KZ behavior in a topological phase transition with a quadratic band crossing. We used dual perspectives: momentum space in clean systems and real space in disordered systems. Both perspectives give consistent scalings, which are distinct from previously studied linear band crossing models. Further, we go beyond quadratic to higher orders. They also follow KZ behavior but acquire different critical exponents.
Our study extends the KZ behavior in topological transitions beyond the linear band crossing and confirms the consistency between correlation length from Berry curvature and LCM. For future studies, we get some hint from this work that the LCM correlation length may characterize the Anderson insulator transition (crossover) if the disorder is strong.
@@ Appendix A Extraction of correlation length from the local Chern marker
Figure 8: (a). The distribution of the LCM in ground state. (b). Disorder-averaged autocorrelation function calculated from the LCM. Here, the NNN hopping strength is set to $J^{\prime}=1$, and the average is performed over $20$ disorder realizations $\delta V=0.05$ and size $N=50$.
Fig.8(a) shows the real-space distribution of the local Chern marker (LCM) in ground state. To estimate the size of inhomogenous ”domains” in the LCM, we calculated the normalized autocorrelation function of the real-space distribution function $\delta c(r)$:
@@
where $\delta c(\mathbf{r})=c(\mathbf{r})-\bar{c}$ is the LCM fluctuation. Similar to previous work, we determine the correlation length in the LCM as the distance at which the disorder-averaged autocorrelation function crosses zero Ulčakar et al. (2020); Sun et al. (2022), $\bar{R}(\xi_{r})=0$. The disorder-averaged autocorrelation function of the LCM in ground state, is shown in Fig.8(b).
Additionally, there is an alternative method to calculate correlation length. Another option is to use the nonlocal Chern marker, which is the off-diagonal elements of the Chern marker operator in Eq.4. This method has a direct relationship with the Fourier transform of the Berry curvatureMolignini et al. (2023).
@@ Appendix B Alternative characterization of the freeze-out time in momentum space
Figure 9: (a). The dynamics of correlation length in momentum space in a rescaled coordinate with $J^{\prime}=1$. Curves collapse near the transition point. The freeze-out time $t_{\mathrm{f}}$ can be determined by the first peak after the transition. The inset shows the same data in a non-rescaled coordinate. (b). Freeze-out time $t_{\mathrm{f}}$ has a power-law scaling with the quench time $t_{\mathrm{f}}\sim\tau^{1/2}$. (c). Correlation length at freeze-out time $t_{\mathrm{f}}$ shows a scaling behavior $\xi_{k}(t_{\mathrm{f}})\sim\tau^{1/4}$.
In the main text, we use the nonequilibrium Hall response to determine the freeze-out time during the quench in clean systems. The freeze-out time by definition characterizes the time when the system’s relaxation rate matches the cooling rate. That is to say, it could be seen as a time delay when the system realizes the transition has occurred. So, the time delay of the trend of divergence of the correlation length can also be utilized to identify the freeze-out time $t_{\mathrm{f}}$. This method can be regarded as a supplement to the Hall response criterion and can be generalized to real space for disordered systems directly.
@@
Appendix C Alternative characterization of the freeze-out time in real space
Figure 10: (a). The dynamics of the spatially averaged Chern marker $\mathcal{C}$ in a rescaled coordinate with size $N=50$. Curves show oscillation after the transition. The freeze-out time $t_{\mathrm{f}}$ can be determined by the first dip after the transition. The inset shows the same data in a non-rescaled coordinate. (b). Freeze-out time $t_{\mathrm{f}}$ has a power-law scaling with the quench time $t_{\mathrm{f}}\sim\tau^{1/2}$. (c). Correlation length at freeze-out time $t_{\mathrm{f}}$ shows a scaling behavior $\xi_{r}(t_{\mathrm{f}})\sim\tau^{1/4}$.
In the main text, we determine the freeze-out time in disordered systems during the quench by the time delay at which the real-space correlation length reaches its maximum after the transition. For disordered systems, we find that the freeze-out time during the quench can also be determined by the time of the first dip in the spatially averaged Chern marker $\mathcal{C}$ after the transition.
Spatially averaged Chern marker $\mathcal{C}$ is determined by the site-averaged and disorder-averaged LCM. It is quantized and equal to clean-band Chern number of the ground state in thermodynamic limit. But during the quench dynamics, $\mathcal{C}$ shows continuous oscillation after the transition point, and finally relaxes to the ground-state value, just similar as the Hall response (not shown within the time window of Fig.10, because the relaxation is quite slow and may take time even longer than $\tau$). So, we can choose the first dip of the oscillation as a signature of the freeze-out time $t_{\mathrm{f}}$. Fig.10 shows the extracted freeze-out time and corresponding unfrozen correlation lengths. It shows the scaling of the freeze-out time $t_{\mathrm{f}}$ to the quench time $\tau$. And the unfrozen correlation length $\xi_{k}(t_{\mathrm{f}})$ also follows KZ’s prediction.